# Noise-Aware Entanglement Generation Protocols for Superconducting Qubits with impedance-matched FBAR Transducers

### v2

@@ Abstract
Connecting superconducting quantum processors to telecommunications-wavelength quantum networks is critically necessary to enable distributed quantum computing, secure communications, and other applications. Optically-mediated entanglement heralding protocols offer a near-term solution that can succeed with imperfect components, including sub-unity efficiency microwave-optical quantum transducers. The viability and performance of these protocols relies heavily on the properties of the transducers used: the conversion efficiency, resonator lifetimes, and added noise in the transducer directly influence the achievable entanglement generation rate and fidelity of an entanglement generation protocol. Here, we use an extended Butterworth-van Dyke (BVD) model to optimize the conversion efficiency and added noise of a Thin Film Bulk Acoustic Resonator (FBAR) piezo-optomechanical transducer. We use the outputs from this model to calculate the fidelity of one-photon and two-photon entanglement heralding protocols in a variety of operating regimes. For transducers with matching circuits designed to either minimize the added noise or maximize conversion efficiency, we theoretically estimate that entanglement generation rates of greater than $160\;\mathrm{kHz}$ can be achieved at moderate pump powers with fidelities of $>90\%$. This is the first time a BVD equivalent circuit model is used to both optimize the performance of an FBAR transducer and to directly inform the design and implementation of an entanglement generation protocol. These results can be applied in the near term to realize quantum networks of superconducting qubits with realistic experimental parameters.
preprint: Distribution D
@@ I Introduction
Figure 1: (a) Schematic of the constituent modes of a piezo-optomechanical transducer and their interactions with each other and the outside world for (top) a single-ring device and (bottom) a double-ring device. $g_{EM}$ is the electromechanical coupling rate, $g_{OM}$ is the optomechanical coupling rate, $\eta_{e}$ is the microwave signal coupling efficiency, $\kappa_{e}$ is the microwave mode ($\hat{a}_{e}$) loss $J$ is the optical-optical coupling rate, rate, $\gamma_{i}$ is the mechanical mode ($\hat{b}_{m}$) intrinsic linewidth, $\kappa_{i}$ is the optical mode ($\hat{a}_{o}$, $\hat{a}_{A}$, $\hat{a}_{S}$ ) intrinsic linewidth, and $\eta_{o}$ is the optical mode external coupling efficiency. (b) Side profile of an example material stack for an FBAR transducer, adapted from [26]. The dashed line denotes the axis of rotational symmetry. General estimates of layer thicknesses are shown, and should be understood as rough guides. Specific FBAR devices may have slightly modified dimensions. (c) Thévenin equivalent circuit for a piezo-optomechanical transducer, including a matching network. $Z_{tx}$ is the input impedance, $L$ is the matching inductor, $C_{T}$ is the matching capacitor, $R_{L}$ is the matching inductor’s resistance, $C_{0}$ is the static capacitance, $R_{0}$ is the static resistance, $R_{m},L_{m},C_{m}$ are the motional resistance, inductance, and capacitance, respectively, $R_{OM\pm}$ are the equivalent resistances for the optical signal and noise outputs, and $V_{tx},V_{L},V_{m},V_{o}$ are the Thévenin equivalent voltage sources which accompany every resistive element. for the Adapted from [27].
@@ I.1 The FBAR Piezo-optomechanical transducer
Transducers based on optomechanical crystal (OMC) cavities have achieved high conversion efficiencies[25, 32, 33]. However, these devices suffer from high added noise quanta per transduced photon, as the OMC cavity is susceptible to optical absorption-based heating [34], leading to non-negligible thermal Brownian motion of the mechanical modes [24]. Recently, two-dimensional OMCs have been developed to enhance thermalization [35, 36] and have achieved thermal occupancies of the phonon cavity $<1$ at high optical pump repetition rates with optical cavity occupations on the order of $10^{3}$ photons. These results motivate the exploration of bulk acoustic resonators for even better thermalization, promising acoustic cavity ground state operation at even higher optical cavity occupations.
@@
$H_{int}=\\ \hbar\sum_{i=e,o}g_{iM}(\hat{b}_{m}e^{-i\omega_{m}t}+\hat{b}^{{\dagger}}_{m}e^{i\omega_{m}t})(\hat{a}_{i}e^{-i\Delta_{i}t}+\hat{a}^{{\dagger}}_{i}e^{i\Delta_{i}t})\$ (1)
@@
with pump-optical mode detuning $\Delta_{o}$. We can select one of two possible optomechanical interactions by red-detuning our pump laser pulse such that $\Delta_{o}=\omega_{m}$ or blue-detuning the pulse such that $\Delta_{o}=-\omega_{m}$ [2]. Multiplying out the optomechanical interaction terms and discarding high-frequency terms oscillating at ±$\pm 2\omega_{m}$, we have, for the red-detuned case,
@@
yielding a beamsplitter-like interaction between the optical cavity mode and the mechanical mode, where excitations are swapped between modes. The red-detuned optical pump, together with a resonant microwave pump, enables direct microwave-to-optical transduction. A blue-detuned pump $\Delta_{o}=\omega_{m}$ leads to optomechanical two-mode squeezing interaction:
@@ I.2 Butterworth-van Dyke model and impedance matching
Figure 2: Diagram of an entanglement heralding setup between two nominally identical quantum network nodes consisting of generic superconducting circuits and microwave-optical transducers, which are located within cryostats operating at $20\;\mathrm{mK}$ base temperature. A single-ring transducer is depicted. A laser sends red-detuned pump pulses to the transducers. Top left shows simplified density of states diagrams for red-detuned pumping of single-ring and double-ring FBAR transducers. Microwave electronics send drive pulses to the superconducting qubits via coaxial cables (not shown) to initialize them in the excited state. Photons output from each transducer coupled off-chip via cryogenic fiber bonds, with a typical loss of $3-10\;\mathrm{dB}$, then undergo pump filtering, also with a typical loss of $3-10\;\mathrm{dB}$, before impinging on a Bell state analyzer (BSA) with a 50:50 beamsplitter. For the Type I protocol, single photon detection events in either photodetector herald entanglement between the remote superconducting qubits. For the Type II protocol, coincident detection events (coins) do the same, using either a balanced or unbalanced BSA.
@@ I.3 Entanglement heralding for superconducting qubits
Often referred to as the DCLZ [9] or Barrett-Kok [10] protocol, optically-mediated entanglement of remote matter-based qubits involves interfering flying photons emitted by those qubits in a Bell state analyzer. The beamsplitter within the analyzer erases which-path information before the photon(s) are incident upon single photon detectors. This information erasure allows one to project entanglement onto the remote quantum memories. These protocols can broadly be classified as one-photon (Type I) or two-photon (Type II) protocols [48].
A Type-I protocol is that which uses an optical single-photon state to establish a distributed Bell state between matter-based qubits at remote network nodes. It will only succeed when one photon is emitted, in total, from both network nodes. As detailed below, Type-I protocols are susceptible to sources of infidelity from two-photon emission events. They are also sensitive to path-length variations within the experimental setup, where even slight changes in the length of the network links can change the relative phase of the entangled state, leading to decoherence.
@@
When we consider heralding protocols in the context of superconducting qubits and transducers, even more complexities are introduced. As discussed above, the transducer can operate in different regimes depending on the frequency of the optical pump with respect to the optical cavity mode(s). Spontaneous parametric down-conversion (SPDC) operation is induced by an optical pump blue-detuned from the optical cavity (Equation LABEL:eq:hint_blue ),2 whereas a red-detuned optical pump yields a beam-splitter-type interaction that can be utilized for quantum coherent state transfer (direct transduction, Equation LABEL:eq:hint_red) [50]. SPDC operation generates entangled microwave-optical photon pairs within the transducer, where the microwave photon can then be swapped into a superconducting qubit. Zhong et al.[29] proposed a Type-II heralding protocol utilizing blue-detuned pumping for SPDC operation of the transducer and time-bin encoded optical photons. Recently, non-classically correlated microwave-telecom photon pair generation has been experimentally demonstrated [51, 32].
@@ II Impedance matching networks for FBAR transducers
Figure 3: A modified equivalent circuit for a transducer with two optical ring resonators. The three sub-circuits are the matching network (blue), mechanical sub-circuit (green) and the optical sub-circuit (orange). As compared to the equivalent circuit shown in Figure 1(c), the optical sub-circuit has been changed here. The optomechanical coupling is now represented by the effective resistance $R_{OM}$ and the effective resistors $R_{OO\pm}$, representing the upper and lower optical sidebands, have been added in parallel with $R_{OM}$.
@@ II.1 The BVD model
One might notice that for our previous condition $C_{m}\ll C_{0}$ we find $k_{eff}\ll 1$ and $\omega_{s}\approx\omega_{p}$. The resonant frequency relates to the series and parallel frequencies by $\omega_{m}=(\omega_{s}+\omega_{p})/2$, and therefore a reasonable approximation exists as $\omega_{m}\approx\omega_{s}.$ This approximation will prove to be beneficial in the next section. Far away from resonance, the device behaves as a capacitor, so its capacitance can be extracted from the admittance curve using the relation $Im(Y_{\omega})\rightarrow\omega C_{0}$. As such, running a finite-element simulation of a device and plotting its admittance curve allows us to extract $k_{eff}^{2}$ and $C_{0}$ for our specific device geometry and inform the equivalent circuit model. Moving beyond the standard BVD model, the modified BVD (mBVD) model includes an additional static resistance $R_{0}$ to the static arm of the piezo-mechanical device, which captures the dielectric loss due to energy dissipation in the piezoelectric layer. The mBVD model has been found to more accurately model the behavior of FBAR devices[8], so we use it here. See Appendix A for more details.
@@ II.2 The Extended BVD model for optomechanical transducers
Before moving on, it’s important to clarify an assumption that is made in this work and in previous work on this model [46, 57] : we are assuming the electrical and optical modes are strongly coupled to a single acoustic mode, and weakly coupled to all other acoustic modes. In practice this may not be the case, particularly in unreleased FBAR transducers with a dense forest of acoustic modes with a free spectral range of $<20\;\mathrm{MHz}$ [37]. Generalizing the equivalent electrical circuits for transducers with multiple acoustic modes coupled in parallel is an area of potential future exploration.
@@ II.2.1 The matching circuit
For a superconducting circuit, we can set $R_{L}=0$. This is a fairly accurate assumption that is explored in detail in Appendix B.
@@ II.2.2 The Optomechanical sub-circuit
The optomechanical coupling between this optical mode and the mechanical mode is represented by frequency-independent effective resistances $R_{OM,\pm}$. The positive resistance $R_{OM,+}$ represents the desired output mode (upper sideband for direct transduction), and the negative resistance $-R_{OM,-}$ represents the optical noise channel due to the Stokes process creating photons in the lower sideband, which excite spurious phonons in the mechanical mode, creating noise photons in the upper sideband [27].
@@
Here $\mathcal{C}_{OM}=\frac{4g_{OM}^{2}}{\gamma_{m}\kappa_{o}}$ is the optomechanical cooperativity and the Lorentzian sideband amplitudes ±$\mathscr{L}^{2}_{\pm}$ are
±$\mathscr{L}^{2}_{\pm}=\frac{(\kappa_{o}/2)^{2}}{(\kappa_{o}/2)^{2}+(\omega_{m}\pm\Delta)^{2}}$ (22)
@@ II.2.3 Figures of Merit
From the equivalent circuit we can also extract expressions for added noise during transduction, due to the presence of both optical amplification noise and thermo-mechanical noise. Optical amplification noise, or Raman noise, occurs when a pump photon produces an output photon at the angular frequency $\omega_{pump}-\omega_{m}$ via a Stokes process. From two-mode squeezing, this process also produces a phonon which is then transduced into the upper sideband as a noise photon.
@@
Details of how we modified the extended BVD model to include the static resistance $R_{0}$ can be found in Appendix A. Now that we have defined the figures of merit of our transducer as they relate to our model, we can proceed to design an impedance matching network.
@@ II.2.4 Designing the matching network
Detailed derivations behind this brief summary section can be found in reference [27]. In Sec. II.4 below, we use this formalism to design impedance matching networks for FBAR transducers. Before detailing these results, we must ask, can the formalism be modified to include an additional cooperativity? If so, how? Such a modification is necessary for transducers that utilize two coupled optical cavities. We explore this question next.
@@
[…]
Electromechanical coupling factor $k_{eff}^{2}$ $34.3E-$ 3
Optical Cavity Photon Number $n_{cav}$ $91E6-2.5E$ 9
[…]

### v3

@@ Abstract
Connecting superconducting quantum processors to telecommunications-wavelength quantum networks is critically necessary to enable distributed quantum computing, secure communications, and other applications. Optically-mediated entanglement heralding protocols offer a near-term solution that can succeed with imperfect components, including sub-unity efficiency microwave-optical quantum transducers. The viability and performance of these protocols relies heavily on the properties of the transducers used: the conversion efficiency, resonator lifetimes, and added noise in the transducer directly influence the achievable entanglement generation rate and fidelity of an entanglement generation protocol. Here, we use an extended Butterworth-van Dyke (BVD) model to optimize the conversion efficiency and added noise of a Thin Film Bulk Acoustic Resonator (FBAR) piezo-optomechanical transducer. We use the outputs from this model to calculate the fidelity of one-photon and two-photon entanglement heralding protocols in a variety of operating regimes. For transducers with matching circuits designed to either minimize the added noise or maximize conversion efficiency, we theoretically estimate that entanglement generation rates of greater than $160\;\mathrm{kHz}$ can be achieved at moderate pump powers with fidelities of $90\%$. This is the first time a BVD equivalent circuit model is used to both optimize the performance of an FBAR transducer and to directly inform the design and implementation of an entanglement generation protocol. These results can be applied in the near term to realize quantum networks of superconducting qubits with realistic experimental parameters.
@@ I Introduction
Figure 1: (a) Schematic of the constituent modes of a piezo-optomechanical transducer and their interactions with each other and the outside world for (top) a single-ring device and (bottom) a double-ring device. $g_{EM}$ is the electromechanical coupling rate, $g_{OM}$ is the optomechanical coupling rate, $\eta_{e}$ is the microwave signal coupling efficiency, $\kappa_{e}$ is the microwave mode ($\hat{a}_{e}$) loss $J$ is the optical-optical coupling rate, rate, $\gamma_{i}$ is the mechanical mode ($\hat{b}_{m}$) intrinsic linewidth, $\kappa_{i}$ is the optical mode ($\hat{a}_{o}$, $\hat{a}_{A}$, $\hat{a}_{S}$ ) intrinsic linewidth, and $\eta_{o}$ is the optical mode external coupling efficiency. (b) Side profile of an example material stack for an FBAR transducer, see[26]. The dashed line denotes the axis of rotational symmetry. General estimates of layer thicknesses are shown, and should be understood as rough guides. Specific FBAR devices may have slightly modified dimensions. (c) Thévenin equivalent circuit for a piezo-optomechanical transducer, including a matching network. $Z_{tx}$ is the input impedance, $L$ is the matching inductor, $C_{T}$ is the matching capacitor, $R_{L}$ is the matching inductor’s resistance, $C_{0}$ is the static capacitance, $R_{0}$ is the static resistance, $R_{m},L_{m},C_{m}$ are the motional resistance, inductance, and capacitance, respectively, $R_{OM\pm}$ are the equivalent resistances for the optical signal and noise outputs, and $V_{tx},V_{L},V_{m},V_{o}$ are the Thévenin equivalent voltage sources which accompany every resistive element. See [27].
@@ I.1 The FBAR Piezo-optomechanical transducer
Transducers based on optomechanical crystal (OMC) cavities have achieved high conversion efficiencies[25, 32, 33]. However, these devices suffer from high added noise quanta per transduced photon, as the OMC cavity is susceptible to optical absorption-based heating [34], leading to non-negligible thermal Brownian motion of the mechanical modes [24]. Recently, two-dimensional OMCs have been developed to enhance thermalization [35, 36] and have achieved thermal occupancies of the phonon cavity <1<1 at high optical pump repetition rates with optical cavity occupations on the order of $10^{3}$ photons. These results motivate the exploration of bulk acoustic resonators for even better thermalization, promising acoustic cavity ground state operation at even higher optical cavity occupations.
@@
$H_{int}=\\ \hbar\sum_{i=e,o}g_{iM}(\hat{b}_{m}e^{-i\omega_{m}t}+\hat{b}^{{\dagger}}_{m}e^{i\omega_{m}t})(\hat{a}_{i}e^{-i\Delta_{i}t}+\hat{a}^{{\dagger}}_{i}e^{i\Delta_{i}t})\\$ (1)
@@
with pump-optical mode detuning $\Delta_{o}$. We can select one of two possible optomechanical interactions by red-detuning our pump laser pulse such that $\Delta_{o}=\omega_{m}$ or blue-detuning the pulse such that $\Delta_{o}=-\omega_{m}$ [2]. Multiplying out the optomechanical interaction terms and discarding high-frequency terms oscillating at $\pm 2\omega_{m}$, we have, for the red-detuned case,
@@
yielding a beamsplitter-like interaction between the optical cavity mode and the mechanical mode, where excitations are swapped between modes. The red-detuned optical pump, together with a resonant microwave pump, enables direct microwave-to-optical transduction. A blue-detuned pump $\Delta_{o}=\omega_{m}$ leads to an optomechanical two-mode squeezing interaction:
@@ I.2 Butterworth-van Dyke model and impedance matching
Figure 2: Diagram of an entanglement heralding setup between two nominally identical quantum network nodes with superconducting circuits and microwave-optical transducers, which are located within cryostats. A single-ring transducer is depicted. A laser sends red-detuned pump pulses to the transducers, and qubits are initialized to their excited state. Photons emitted by each transducer coupled off-chip via cryogenic fiber bonds, then undergo pump filtering before impinging on a 50:50 beamsplitter. Typical loss values are labeled. Top left: density of states diagrams for red-detuned pumping of single-ring and double-ring FBAR transducers Top right: for the Type I protocol, single photon clicks in either detector herald entanglement between the remote superconducting qubits. For the Type II protocol, coincident detection events (coins.) do the same, using a balanced or unbalanced interferometer.
@@ I.3 Entanglement heralding for superconducting qubits
Often referred to as the DLCZ [9] or Barrett-Kok [10] protocol, optically-mediated entanglement of remote matter-based qubits involves interfering flying photons emitted by those qubits in a Bell state analyzer. The beamsplitter within the analyzer erases which-path information before the photon(s) are incident upon single photon detectors. This information erasure allows one to project entanglement onto the remote quantum memories. These protocols can broadly be classified as one-photon (Type I) or two-photon (Type II) protocols [49].
A Type-I protocol is that which uses an optical single-photon state to establish a distributed Bell state between matter-based qubits at remote network nodes. It will only succeed when one photon is emitted, in total, from both network nodes. As detailed below, Type-I protocols are susceptible to sources of infidelity from two-photon emission events. They are also sensitive to path-length variations within the experimental setup, where even slight changes in the length of the network links can change the relative phase of the entangled state, leading to decoherence. In contrast, Type-II protocols use two optical photons, one emitted from each of two nodes, and measure coincident detection events between them.
@@
When we consider heralding protocols in the context of superconducting qubits and transducers, even more complexities are introduced. As discussed above, the transducer can operate in different regimes depending on the frequency of the optical pump with respect to the optical cavity mode(s). Spontaneous parametric down-conversion (SPDC) operation is induced by an optical pump blue-detuned from the optical cavity (Equation LABEL:eq:hint_blue ), whereas a red-detuned optical pump yields a beam-splitter-type interaction that can be utilized for quantum coherent state transfer (direct transduction, Equation LABEL:eq:hint_red) [51]. SPDC operation generates entangled microwave-optical photon pairs within the transducer, where the microwave photon can then be swapped into a superconducting qubit. Zhong et al.[29] proposed a Type-II heralding protocol utilizing blue-detuned pumping for SPDC operation of the transducer and time-bin encoded optical photons. Recently, non-classically correlated microwave-telecom photon pair generation has been experimentally demonstrated [52, 32].
@@
While previous circuit-based treatments of FBAR transducers focus on device characterization[26, 46, 27], our approach is the first to directly connect BVD circuit parameters to quantum entanglement protocol performance metrics. This connection allows us to concurrently optimize both conversion efficiency and added noise for entanglement applications, whereas previous work optimized these figures of merit separately.
Our treatment also generalizes the extended BVD model in several key ways: (a) including two optical resonators to model photonic molecule architectures, whose optical emission frequency can be tuned via the symmetric/antisymmetric supermode splitting, (b) including parasitic circuit elements and their compensation in matching network design, and (c) connecting device-level optimization directly to protocol-level performance. Unlike previous theoretical explorations of entanglement heralding protocols [49, 12], we provide a complete device engineering framework that bridges transducer physics to quantum networking applications.
@@ II Impedance matching networks for FBAR transducers
Figure 3: A modified equivalent circuit for a transducer with two optical ring resonators. The three sub-circuits are the matching network (blue), mechanical sub-circuit (green) and the optical sub-circuit (red). As compared to the equivalent circuit shown in Figure 1(c), the optical sub-circuit has been changed here. The optomechanical coupling is now represented by the effective resistance $R_{OM}$ and the effective resistors $R_{OO\pm}$, representing the upper and lower optical sidebands, have been added in parallel with $R_{OM}$.
@@ II.1 The BVD model
One might notice that for our previous condition $C_{m}\ll C_{0}$ we find $k_{eff}\ll 1$ and $\omega_{s}\approx\omega_{p}$. The resonant frequency relates to the series and parallel frequencies by $\omega_{m}=(\omega_{s}+\omega_{p})/2$, and therefore a reasonable approximation exists as $\omega_{m}\approx\omega_{s}.$ This approximation will prove to be beneficial in the next section. Far away from resonance, the device behaves as a capacitor, so its capacitance can be extracted from the admittance curve using the relation $Im(Y_{\omega})\rightarrow\omega C_{0}$. As such, running a finite-element simulation of a device and plotting its admittance curve allows us to extract $k_{eff}^{2}$ and $C_{0}$ for our specific device geometry and inform the equivalent circuit model. Moving beyond the standard BVD model, the modified BVD (mBVD) model includes an additional static resistance $R_{0}$ to the static arm of the piezo-mechanical device, which captures the dielectric loss due to energy dissipation in the piezoelectric layer. The mBVD model has been found to more accurately model the behavior of FBAR devices[8], so we use it here. See Appendix B for more details.
@@ II.2 The Extended BVD model for optomechanical transducers
Before moving on, it’s important to clarify an assumption that is made in this work and in previous work on this model [47, 58] : we are assuming the electrical and optical modes are strongly coupled to a single acoustic mode, and weakly coupled to all other acoustic modes. In practice this may not be the case, particularly in unreleased FBAR transducers with a dense forest of acoustic modes with a free spectral range of $20\;\mathrm{MHz}$ [37]. Generalizing the equivalent electrical circuits for transducers with multiple acoustic modes coupled in parallel is an area of potential future exploration.
@@ II.2.1 The matching circuit
For a superconducting circuit, we can set $R_{L}=0$. This is a fairly accurate assumption that is explored in detail in Appendix C.
@@ II.2.2 The Optomechanical sub-circuit
The optomechanical coupling between this optical mode and the mechanical mode is represented by frequency-independent effective resistances $R_{OM,\pm}$. The positive resistance $R_{OM,+}$ represents the desired output mode (upper sideband for direct transduction), and the negative resistance $R_{OM,-}$ represents the optical noise channel due to the Stokes process creating photons in the lower sideband, which excite spurious phonons in the mechanical mode, creating noise photons in the upper sideband [27].
@@
Here $\mathcal{C}_{OM}=\frac{4g_{OM}^{2}}{\gamma_{m}\kappa_{o}}$ is the optomechanical cooperativity and the Lorentzian sideband amplitudes $\mathscr{L}^{2}_{\pm}$ are
$\mathscr{L}^{2}_{\pm}=\frac{(\kappa_{o}/2)^{2}}{(\kappa_{o}/2)^{2}+(\omega_{m}\pm\Delta)^{2}}$ (22)
@@ II.2.3 Figures of Merit
From the equivalent circuit we can also extract expressions for added noise during transduction, due to the presence of both optical amplification noise and thermo-mechanical noise. Optical amplification noise, or Raman noise, occurs when a pump photon produces an output photon at the angular frequency $\omega_{pump}-\omega_{m}$ via a Stokes process. From two-mode squeezing, this process also produces a phonon which is then transduced into the upper sideband as an optical Raman noise photon.
@@
Details of how we modified the extended BVD model to include the static resistance $R_{0}$ can be found in Appendix B. Now that we have defined the figures of merit of our transducer as they relate to our model, we can proceed to design an impedance matching network.
@@ II.2.4 Designing the matching network
Detailed derivations behind this brief summary section can be found in reference [27]. In Sec. II.5 below, we use this formalism to design impedance matching networks for FBAR transducers. Before detailing these results, we must ask, can the formalism be modified to include an additional cooperativity? If so, how? Such a modification is necessary for transducers that utilize two coupled optical cavities. We explore this question next.
@@
optomechanical coupling rate $g_{OM,0}/2\pi$ $400\;\textrm{Hz}$
Optical cavity frequency $\omega_{o}/2\pi$ $193\;\textrm{THz}$
Optical intrinsic linewidth $\kappa_{i}/2\pi$ $25\;\textrm{MHz}$
Optical external coupling rate $\kappa_{ext}/2\pi$ $125\;\textrm{MHz}$
Mechanical mode frequency $\omega_{m}/2\pi$ $3.285\;\textrm{GHz}$
Mechanical intrinsic linewidth $\gamma_{i}/2\pi$ $2.6\;\textrm{MHz}$
Electromechanical coupling factor $k_{eff}^{2}$ $4.3\times 10^{-3}$
Optical Cavity Photon Number $n_{cav}$ $1\times 10^{6}-2.5\times 10^{9}$
Static capacitance $C_{0}$ $200\;\textrm{fF}$